3)You opened a saving account for your son 4 years ago and deposited $500 at that time. Three years ago, you added another $500 to the account. Last year, you deposited an additional $300 into this account. With an interest rate of 5% compounded annually, how much is in the account today?

4)Mike borrowed $150,000 from ABC Bank for 5 years at an interest rate of 12% compounded monthly. How much is his monthly loan payment?
1
Expert's answer
2020-06-29T15:03:29-0400

3) "500*1.05^4+500*1.05^3+300*1.05=1,501.57"

4) "P=150,000*\frac{0.12/12*(1+0.12/12)^{60}}{(1+0.12/12)^{60}-1}=3,336.67"

P-payment monthly